# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

### v1

@@ A.1 More related work
Deep learning in human activity recognition. Over the last decade, HAR has transitioned from hand-crafted feature extraction to deep learning models capable of automatic feature learning. Early work by Kwapisz et al. (2011) utilized machine learning techniques, such as decision trees and MLPs, to classify activities using features extracted from wearable sensor data. Later, Haresamudram et al. (2019) demonstrated that optimized feature extraction within the Activity Recognition Chain (ARC) could rival or outperform end-to-end deep learning models. Deep learning models, particularly CNNs and LSTMs, have since become dominant in HAR. Bevilacqua et al. (2019) developed a CNN-based model for HAR, while Ha & Choi (2016) introduced CNN-pf and CNN-pff architectures that apply partial and full weight sharing for better feature extraction. Other notable works include Perception-Net Kasnesis et al. (2019), which leverages 2D convolutions for multimodal sensor data, and InnoHAR (Xu et al., 2019), which combines Inception CNN and GRUs for multiscale temporal feature learning. A dual-stream network utilizing convolutional layers and LSTM units, known as ConvLSTM, was employed by Yuki et al. (2018) to analyze complex temporal hierarchies with streams handling different time lengths. The combination of attention mechanisms with recurrent networks to enhance the computation of weights for hidden state outputs has also been demonstrated by DeepConvLSTM Kasnesis et al. (2019) in capturing spatial-temporal features.
@@
Large Language Models for Time-Series Forecasting. LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting. Xue & Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings. Gruver et al. (2023) explores encoding time-series as numerical strings, allowing LLMs like GPT-3 and LLaMA-2 to perform zero-shot forecasting, matching or surpassing the performance of specialized models, while highlighting challenges in uncertainty calibration due to model modifications like RLHF. Zhou et al. (2023a) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks. Jin et al. (2024b) highlights the transformative potential of LLMs for time-series analysis by integrating language models with traditional analytical methods. Jin et al. (2024a) introduces a reprogramming framework that aligns time-series data with natural language processing capabilities, enabling LLMs to perform time-series forecasting without altering the core model structure. Cao et al. (2024) presents TEMPO, a generative transformer framework based on prompt tuning, which adapts pre-trained models for time-series forecasting by decomposing trends, seasonality, and residual information. Sun et al. (2024a) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications. Chang et al. (2024) develops a framework that enhances pre-trained LLMs for multivariate time-series forecasting through a two-stage fine-tuning process and a novel multi-scale temporal aggregation method, outperforming traditional models in both full-shot and few-shot scenarios. Finally, Liu et al. (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
@@
LLMs for Human Activity Recognition. While LLMs like ChatGPT have demonstrated remarkable performance in various NLP tasks, their effectiveness in HAR remains limited due to challenges in interpreting sensor data. These models often struggle to distinguish between activities that share similar objects, requiring more advanced prompt engineering to highlight activity-specific details. Xia et al. (2023) proposed an unsupervised approach to HAR using ChatGPT, leveraging two-stage prompts to infer activities from object sequences without manual descriptions. The method demonstrates superior performance on three benchmark datasets, marking a significant advancement in applying language models to activity recognition tasks. Similarly, Ji et al. (2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
@@ A.2 Data Generation
In this section, we outline the process for generating text data from sensor readings. Our approach utilizes a set of predefined sentence templates, which are randomly selected to create diverse question-answer (QA) pairs. To enhance the variability, we use GPT-4o to generate synonymous versions of the sentences within the templates. Each sentence contains placeholders for relevant numerical values (e.g., timestamps, sensor readings) or textual information. We also choose a variety of synonym expressions for each trend. These values are then inserted into the placeholders to form coherent QA pairs corresponding to the sensor data inputs.
Table 4: Question templates examples.
@@
Description: Kindly provide a detailed analysis of the trend changes observed in the {data}. Please offer a comprehensive description of how the trends in the {data} have evolved. I would appreciate a thorough explanation of the trend fluctuations that occurred within the {data}. Could you please examine the {data} in depth and explain the trend shifts observed step by step? Detail the {data}’s trend transitions. Could you kindly assess the {data} and provide a description of the trend transformations that took place step by step? Could you analyze the trends observed in the {data} over the specified period step by step? Can you dissect the {data} and explain the trend changes in a detailed manner? What trend changes can be seen in the {data}?
@@
Summary: Could you provide a summary of the main features of the input {data} and the distribution of the trends? Please give an overview of the essential attributes of the input {data} and the spread of the trends. Describe the salient features and trend distribution within the {data}. Give a summary of the {data}’s main elements and trend apportionment. Summarize the {data}’s core features and trend dissemination. Outline the principal aspects and trend allocation of the {data}. Summarize the key features and trend distribution of the {data}. I need a summary of {data}’s main elements and their trend distributions.
Table 5: Answer templates examples.
Description: {start_time}s to {end_time}s: {trend} {start_time} seconds to {end_time} seconds: {trend} {start_time} to {end_time} seconds: {trend} {start_time}-{end_time} seconds: {trend} {start_time}-{end_time}s: {trend} {start_time}s-{end_time}s: {trend}
Summary 1: Number of {trend} trends: {num} Count of {trend} trends: {num} Number of {trend} segments: {num} Count of {trend} segments: {num}
@@
Summary 2: The given {data_name} representing the {sensor_name} sensor readings from {start_time}s to {end_time}s. The {data_name} represents readings taken from the {sensor_name} sensor between {start_time} and {end_time} seconds. The {sensor_name} sensor readings recorded within the {start_time} to {end_time} second timeframe are presented in this {data_name}.
@@
Summary 3: The data exhibits {trend_num} distinct trends, with a total of {change_num} changes in trend observed. Across {trend_num} trends, the data shows {change_num} occurrences of trend shifts. {trend_num} trends are present in the data, with {change_num} instances of trend changes.
@@
Summary 4: To sum up, the data exhibited a {trend_type} trend for a cumulative period of {total_time} seconds. Overall, the data showed a {trend_type} trend over {total_time} seconds. To conclude, the trend was {trend_type} over a period of {total_time} seconds.
Summary 5: The overall trend is {overall_trend}. The primary trend detected is {overall_trend}. Looking at the big picture, the trend is {overall_trend}.
@@
The system prompt guides the model on how to respond to the questions generated, incorporating characteristics specific to each dataset, such as sensor frequency and other dataset-specific attributes. These tailored prompts ensure the responses are aligned with the unique properties of the sensor data. Below is the template of the system prompts (instructions) used for each dataset:
@@
  • A dialogue between a curious researcher and an AI assistant. The AI analyzes a sensor time-series dataset (N points, {sample_rate}Hz sampling rate) to answer specific questions. This interaction demonstrates the AI’s data analysis skills and the potential of human-AI collaboration in interpreting complex data.
A.3 GPT-4o Prompt for Sensor Data Trend Analysis
Table 6 outlines the system prompt used to generate trend-descriptive texts from sensor data, providing a structured format for GPT-4o to analyze and respond to specific questions about the data. This prompt serves as a standardized way for GPT-4o to examine the time-series data and produce outputs that can then be directly compared to the descriptions generated by our SensorLLM model. By using this prompt, we aim to evaluate GPT-4o’s ability to interpret numerical sensor data and generate accurate trend descriptions. The responses generated by GPT-4o are assessed against both human evaluations and NLP metrics to benchmark its performance against SensorLLM, allowing us to critically analyze the differences in how both models understand and interpret time-series data trends. This comparison helps demonstrate the effectiveness of our Sensor-Language Alignment Stage of SensorLLM.
@@ A.3 GPT-4o Prompt for Sensor Data Trend Analysis
A.4 Chronos
Chronos (Ansari et al., 2024) is an innovative framework for pretrained probabilistic time series models. It tokenizes time series data into a fixed vocabulary through scaling and quantization, then trains T5-based (Raffel et al., 2020) language models on these tokenized sequences using cross-entropy loss. Pretrained on a diverse corpus of public and synthetic datasets, Chronos models demonstrate remarkable performance, outperforming other methods on familiar datasets and showing competitive zero-shot capabilities on unseen tasks. This approach leverages data from various domains to enhance forecasting accuracy, potentially simplifying time series analysis pipelines across different fields.
Time-series Tokenization and Quantization. Chronos tokenizes real-valued time-series data through a two-step process: normalization and quantization. First, mean scaling is applied for normalization, where each observation is divided by the mean of its absolute values:
$\tilde{x}=\frac{x}{\text{mean}(\lvert x\rvert)}.$
This ensures that the data is scaled consistently across different series. After normalization, the values are quantized into discrete tokens. Quantization is performed by selecting $B$ bin centers $c_{1}<\dots<c_{B}$ along with $B-1$ edges $b_{i}$ separating them. Each real value is mapped to a token based on the quantization function $q(x)$, defined as:
$q(x)=\begin{cases}1&\text{if}\ -\infty\leq x<b_{1},\\ 2&\text{if}\ b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if}\ b_{B-1}\leq x<\infty,\end{cases}$
where each token represents a quantized range of values. This quantization ensures that the continuous time-series data can be represented as discrete tokens suitable for language models.
To further facilitate the modeling, special tokens such as PAD and EOS are added to the time-series vocabulary $V_{ts}$ to handle sequence padding and denote the end of sequences. This allows Chronos to efficiently process variable-length sequences and integrate them with language modeling tasks, treating the time-series data as token sequences.
Objective Function. In Chronos, the model uses a categorical distribution over the quantized time-series tokens from the vocabulary $V_{ts}$. The objective is to minimize the cross-entropy between the predicted categorical distribution and the ground-truth labels, including EOS tokens. The loss function for a single time series is defined as:
$\ell(\theta)=-\sum_{h=1}^{H+1}\sum_{i=1}^{|V_{ts}|}{\bf\mathbf{1}}(z_{C+h+1}=i)\log p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$
where $C$ represents historical context of length and $H$ represents a forecast horizon, and $p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$ is the predicted distribution. The loss is averaged over batches of time series.
This categorical output distribution offers two key advantages: (i) it requires no modification to the language model, allowing easy integration with existing architectures; and (ii) it supports learning arbitrary distributions, making it flexible for diverse time-series datasets.
@@ A.5 Evaluation Metrics for Sensor-Language Alignment Stage
BLEU-1 (Papineni et al., 2002). BLEU (Bilingual Evaluation Understudy) is a precision-based metric commonly used to evaluate machine-generated text by comparing it to reference texts. BLEU-1 focuses on unigram (single-word) overlap, assessing the lexical similarity between the generated and reference text. While useful for measuring word-level matches, BLEU-1 does not capture deeper semantic meaning, making it most effective for surface-level alignment.
ROUGE-1 and ROUGE-L (Lin, 2004). ROUGE (Recall-Oriented Understudy for Gisting Evaluation) evaluates the recall-oriented overlap between generated text and reference text. ROUGE-1 focuses on unigram recall, similar to BLEU-1 but emphasizing how much of the reference text is captured. ROUGE-L measures the longest common subsequence, assessing both precision and recall in terms of structure and content overlap, though it does not evaluate semantic accuracy.
METEOR (Banerjee & Lavie, 2005). METEOR (Metric for Evaluation of Translation with Explicit Ordering)combines precision and recall, with additional alignment techniques such as stemming and synonym matching. Unlike BLEU and ROUGE, METEOR accounts for some degree of semantic similarity. However, its emphasis is still on word-level alignment rather than factual accuracy or meaning.
SBERT (Reimers & Gurevych, 2019). SBERT (Sentence-BERT)  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is a metric that generates sentence embeddings using the BERT architecture. It computes cosine similarity between embeddings of the generated and reference texts, providing a deeper assessment of semantic similarity beyond lexical matches.
SimCSE (Gao et al., 2021). SimCSE (Simple Contrastive Sentence Embedding)  [^3] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large introduces a contrastive learning approach to fine-tune language models for sentence embeddings. By applying different dropout masks to the same sentence, it generates positive examples, encouraging similar embeddings for semantically identical sentences while distinguishing different ones.
GPT-4o Evaluation. In addition to the NLP metrics, we also employed GPT-4o as a human-like evaluator. Given its strong reasoning and comprehension abilities, GPT-4o was tasked with scoring the generated text based on its alignment with the ground truth. GPT-4o evaluated the correctness, completeness, and coherence of the trend descriptions and assigned a score from 1 to 5, accompanied by an explanation (see Table 7). This type of evaluation provides insights into how well the generated outputs capture the nuances of sensor data trends in a manner similar to human understanding.
@@
Provide your score (1-5) and a brief explanation in the format: ”score#reason” (e.g., 4#The description of trend changes slightly differs from the ground truth).
@@
Human Evaluation. Finally, five human experts assessed the correctness and quality of the generated trend descriptions. Following the same criteria as GPT-4o, they rated the outputs on a scale from 1 to 5, focusing on the factual accuracy and coherence of the descriptions. This manual evaluation serves as an important benchmark for the model’s performance from a human perspective, ensuring that the generated outputs are not only technically correct but also practically useful for human interpretation.
A.6 Baselines for Task-Aware Tuning Stage

### v2

@@ A.1 More related work
Over the last decade, HAR has transitioned from hand-crafted feature extraction to deep learning models capable of automatic feature learning. Early work by Kwapisz et al. (2011) utilized machine learning techniques, such as decision trees and MLPs, to classify activities using features extracted from wearable sensor data. Later, Haresamudram et al. (2019) demonstrated that optimized feature extraction within the Activity Recognition Chain (ARC) could rival or outperform end-to-end deep learning models. Deep learning models, particularly CNNs and LSTMs, have since become dominant in HAR. Bevilacqua et al. (2019) developed a CNN-based model for HAR, while Ha and Choi (2016) introduced CNN-pf and CNN-pff architectures that apply partial and full weight sharing for better feature extraction. Other notable works include Perception-Net Kasnesis et al. (2019), which leverages 2D convolutions for multimodal sensor data, and InnoHAR (Xu et al., 2019), which combines Inception CNN and GRUs for multiscale temporal feature learning. A dual-stream network utilizing convolutional layers and LSTM units, known as ConvLSTM, was employed by Yuki et al. (2018) to analyze complex temporal hierarchies with streams handling different time lengths. The combination of attention mechanisms with recurrent networks to enhance the computation of weights for hidden state outputs has also been demonstrated by DeepConvLSTM Kasnesis et al. (2019) in capturing spatial-temporal features.
Large Language Models for Time-Series Forecasting.
@@
LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting. Xue and Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings. Gruver et al. (2023) explores encoding time-series as numerical strings, allowing LLMs like GPT-3 and LLaMA-2 to perform zero-shot forecasting, matching or surpassing the performance of specialized models, while highlighting challenges in uncertainty calibration due to model modifications like RLHF. Zhou et al. (2023b) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks. Jin et al. (2024b) highlights the transformative potential of LLMs for time-series analysis by integrating language models with traditional analytical methods. Jin et al. (2024a) introduces a reprogramming framework that aligns time-series data with natural language processing capabilities, enabling LLMs to perform time-series forecasting without altering the core model structure. Cao et al. (2024) presents TEMPO, a generative transformer framework based on prompt tuning, which adapts pre-trained models for time-series forecasting by decomposing trends, seasonality, and residual information. Sun et al. (2024b) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications. Chang et al. (2024) develops a framework that enhances pre-trained LLMs for multivariate time-series forecasting through a two-stage fine-tuning process and a novel multi-scale temporal aggregation method, outperforming traditional models in both full-shot and few-shot scenarios. Finally, Liu et al. (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
LLMs for Human Activity Recognition.
@@
While LLMs like ChatGPT have demonstrated remarkable performance in various NLP tasks, their effectiveness in HAR remains limited due to challenges in interpreting sensor data. These models often struggle to distinguish between activities that share similar objects, requiring more advanced prompt engineering to highlight activity-specific details. Xia et al. (2023) proposed an unsupervised approach to HAR using ChatGPT, leveraging two-stage prompts to infer activities from object sequences without manual descriptions. The method demonstrates superior performance on three benchmark datasets, marking a significant advancement in applying language models to activity recognition tasks. Similarly, Ji et al. (2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
@@ A.2 Data Generation
We generate text data from sensor readings using predefined sentence templates (5, Tables 6,, 7). These templates are randomly selected to create diverse question-answer (QA) pairs. To enhance variability, we employ GPT-4o to generate synonymous variations. Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
Trend Description Templates
{start_time}s to {end_time}s: {trend} {start_time} seconds to {end_time} seconds: {trend} {start_time} to {end_time} seconds: {trend} {start_time}-{end_time} seconds: {trend} {start_time}-{end_time}s: {trend} {start_time}s-{end_time}s: {trend}
Table 5: Examples of answer templates used for trend descriptions.
Trend Description Templates
@@
Kindly provide a detailed analysis of the trend changes observed in the {data}. Please offer a comprehensive description of how the trends in the {data} have evolved. I would appreciate a thorough explanation of the trend fluctuations that occurred within the {data}. Could you examine the {data} in depth and explain the trend shifts observed step by step? Detail the {data}’s trend transitions. Could you assess the {data} and describe the trend transformations step by step? Could you analyze the trends observed in the {data} over the specified period step by step? Can you dissect the {data} and explain the trend changes in a detailed manner? What trend changes can be seen in the {data}?
Summary Templates
@@
Could you provide a summary of the main features of the input {data} and the distribution of the trends? Please give an overview of the essential attributes of the input {data} and the spread of the trends. Describe the salient features and trend distribution within the {data}. Give a summary of the {data}’s main elements and trend apportionment. Summarize the {data}’s core features and trend dissemination. Outline the principal aspects and trend allocation of the {data}. Summarize the key features and trend distribution of the {data}. I need a summary of {data}’s main elements and their trend distributions.
Table 6: Examples of question templates used for trend description and summary generation.
Summary 1: Trend Count
Number of {trend} trends: {num} Count of {trend} trends: {num} Number of {trend} segments: {num} Count of {trend} segments: {num}
Summary 2: Sensor Data Context
@@
The given {data_name} represents {sensor_name} sensor readings from {start_time}s to {end_time}s. The {data_name} contains {sensor_name} sensor readings recorded between {start_time} and {end_time} seconds. The {sensor_name} sensor readings collected from {start_time} to {end_time} seconds are presented in this {data_name}.
Summary 3: Trend Change Statistics
@@
The data exhibits {trend_num} distinct trends, with {change_num} trend changes observed. Across {trend_num} trends, the data shows {change_num} occurrences of trend shifts. {trend_num} trends are present, with {change_num} instances of trend changes.
Summary 4: Cumulative Trend Analysis
@@
To sum up, the data exhibited a {trend_type} trend for a total duration of {total_time} seconds. Overall, the data showed a {trend_type} trend spanning {total_time} seconds. In conclusion, the trend was {trend_type} over {total_time} seconds.
Summary 5: Overall Trend Summary
The overall trend is {overall_trend}. The primary trend detected is {overall_trend}. Looking at the broader pattern, the trend is {overall_trend}.
Table 7: Examples of answer templates used for summaries.
@@
The system prompt instructs the model on how to respond to generated questions, incorporating dataset-specific attributes such as sensor frequency and sampling rate. These tailored prompts ensure responses align with the unique characteristics of each dataset. Below is the system prompt template used for all datasets:
@@
  • A dialogue between a researcher and an AI assistant. The AI analyzes a sensor time-series dataset (N points, sampled at {sample_rate}Hz) to answer specific questions, demonstrating its analytical capabilities and the potential for human-AI collaboration in interpreting sensor data.
A.3 Chronos
Chronos (Ansari et al., 2024) is a pretrained probabilistic time-series framework that tokenizes real-valued time-series data into discrete representations for language model training. It utilizes scaling and quantization to transform time-series data into a fixed vocabulary, enabling T5-based (Raffel et al., 2020) models to learn from tokenized sequences using cross-entropy loss. Pretrained on diverse public and synthetic datasets, Chronos surpasses existing models on familiar datasets and demonstrates strong zero-shot performance on unseen tasks, making it a versatile tool for time-series forecasting across domains.
Time-Series Tokenization and Quantization.
Chronos converts time-series data into discrete tokens through a two-step process: normalization and quantization. Mean scaling is first applied to ensure consistency across different time series:
$\tilde{x}=\frac{x}{\text{mean}(\lvert x\rvert)}$ (5)
Next, the normalized values are quantized using $B$ bin centers $c_{1},\dots,c_{B}$ and corresponding bin edges $b_{1},\dots,b_{B-1}$, mapping real values to discrete tokens via:
$q(x)=\begin{cases}1&\text{if}\ -\infty\leq x<b_{1},\\ 2&\text{if}\ b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if}\ b_{B-1}\leq x<\infty.\end{cases}$ (6)
Special tokens such as PAD and EOS are added to handle sequence padding and denote the end of sequences, allowing Chronos to process variable-length inputs efficiently within language models.
Objective Function.
Chronos models the tokenized time series using a categorical distribution over the vocabulary $V_{ts}$, minimizing the cross-entropy loss:
$\begin{split}\ell(\theta)=-\sum_{h=1}^{H+1}\sum_{i=1}^{|V_{ts}|}\mathbf{1}(z_{C+h+1}=i)\\ \cdot\log p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})\end{split}$ (7)
where $C$ is the historical context length, $H$ is the forecast horizon, and $p_{\theta}$ is the predicted token distribution.
This approach offers two key advantages: (i) Seamless integration with language models, requiring no architectural modifications, and (ii) Flexible distribution learning, enabling robust generalization across diverse time-series datasets.
A.4 GPT-4o Prompt for Sensor Data Trend Analysis
Table 8 presents the system prompt used to generate trend-descriptive texts from sensor data, providing a structured framework for GPT-4o to analyze and respond to specific questions. This standardized prompt ensures consistency in GPT-4o’s interpretation of time-series data, allowing direct comparison with descriptions produced by SensorLLM.
@@ A.4 GPT-4o Prompt for Sensor Data Trend Analysis
We evaluate GPT-4o’s ability to interpret numerical sensor data by assessing its responses against human evaluations and NLP metrics. This comparison benchmarks GPT-4o’s performance against SensorLLM, highlighting differences in how both models process time-series data trends. The results demonstrate the effectiveness of SensorLLM’s Sensor-Language Alignment Stage.
@@ A.5 Evaluation Metrics for Sensor-Language Alignment Stage
BLEU-1 (Papineni et al., 2002).
BLEU (Bilingual Evaluation Understudy) is a precision-based metric commonly used to evaluate machine-generated text by comparing it to reference texts. BLEU-1 focuses on unigram (single-word) overlap, assessing the lexical similarity between the generated and reference text. While useful for measuring word-level matches, BLEU-1 does not capture deeper semantic meaning, making it most effective for surface-level alignment.
ROUGE-1 and ROUGE-L (Lin, 2004).
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) evaluates the recall-oriented overlap between generated text and reference text. ROUGE-1 focuses on unigram recall, similar to BLEU-1 but emphasizing how much of the reference text is captured. ROUGE-L measures the longest common subsequence, assessing both precision and recall in terms of structure and content overlap, though it does not evaluate semantic accuracy.
METEOR (Banerjee and Lavie, 2005).
METEOR (Metric for Evaluation of Translation with Explicit Ordering)combines precision and recall, with additional alignment techniques such as stemming and synonym matching. Unlike BLEU and ROUGE, METEOR accounts for some degree of semantic similarity. However, its emphasis is still on word-level alignment rather than factual accuracy or meaning.
SBERT (Reimers and Gurevych, 2019).
SBERT (Sentence-BERT)  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is a metric that generates sentence embeddings using the BERT architecture. It computes cosine similarity between embeddings of the generated and reference texts, providing a deeper assessment of semantic similarity beyond lexical matches.
SimCSE (Gao et al., 2021).
SimCSE (Simple Contrastive Sentence Embedding)  [^3] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large introduces a contrastive learning approach to fine-tune language models for sentence embeddings. By applying different dropout masks to the same sentence, it generates positive examples, encouraging similar embeddings for semantically identical sentences while distinguishing different ones.
GPT-4o Evaluation.
In addition to the NLP metrics, we also employed GPT-4o as a human-like evaluator. Given its strong reasoning and comprehension abilities, GPT-4o was tasked with scoring the generated text based on its alignment with the ground truth. GPT-4o evaluated the correctness, completeness, and coherence of the trend descriptions and assigned a score from 1 to 5, accompanied by an explanation (see Table 9). This type of evaluation provides insights into how well the generated outputs capture the nuances of sensor data trends in a manner similar to human understanding.
@@
Provide your score (1-5) and a brief explanation in the format: "score#reason" (e.g., 4#The description of trend changes slightly differs from the ground truth).
@@
Human Evaluation.
Finally, five human experts assessed the correctness and quality of the generated trend descriptions. Following the same criteria as GPT-4o, they rated the outputs on a scale from 1 to 5, focusing on the factual accuracy and coherence of the descriptions. This manual evaluation serves as an important benchmark for the model’s performance from a human perspective, ensuring that the generated outputs are not only technically correct but also practically useful for human interpretation.
A.6 Datasets
We used five datasets in our study:
USC Human Activity Dataset (USC-HAD).
USC-HAD (Zhang and Sawchuk, 2012) consists of six sensor readings from body-worn 3-axis accelerometers and gyroscopes, collected from 14 subjects. The data is sampled at 100 Hz across six channels and includes 12 activity class labels. For evaluation, we use data from subjects 13 and 14 as the test set, while the remaining subjects’ data are used for training. A window size $w\in[5,200]$ is used in alignment stage, and $w=200$ with stride of $100$ are used in HAR.
UCI Human Activity Recognition Dataset (UCI-HAR).
UCI-HAR (Anguita et al., 2013) includes data collected from 30 volunteers performing six activities while wearing a smartphone on their waist. The embedded accelerometer and gyroscope sensors sampled data at 50 Hz across six channels. The dataset was partitioned into 70% for training and 30% for testing. A window size $w\in[5,200]$ is used in alignment stage, and $w=128$ with stride of $64$ is used in HAR.
Physical Activity Monitoring Dataset (PAMAP2).
PAMAP2 (Reiss and Stricker, 2012) includes data from nine subjects wearing IMUs on their chest, hands, and ankles. IMUs capture the acceleration, gyroscope, and magnetometer data across 27 channels and include 12 activity class labels. For our experiments, data from subjects 105 and 106 are used as the test set, with the remaining subjects’ data used for training. The sample rate is downsampled from 100 Hz to 50 Hz. A window size $w\in[5,100]$ is used in alignment stage, and $w=100$ with stride of $50$ in HAR.
Mobile Health Dataset (MHealth).
MHealth (Baños et al., 2014) contains body motion and vital sign recordings from ten volunteers. Sensors were placed on the chest, right wrist, and left ankle of each subject. For our experiments, we used acceleration data from the chest, left ankle, and right lower arm, along with gyroscope data from the left ankle and right lower arm, resulting in a total of 15 channels. The data is sampled at 50 Hz and includes 12 activity class labels. Data from subjects 1, 3, and 6 is used as the test set, while the remaining subjects’ data are used for training. We use a window size $w\in[5,100]$ in alignment stage and $w=100$ with stride of $50$ in HAR.
CAPTURE-24.
CAPTURE-24 (Chan et al., 2024) is a large-scale dataset featuring 3-channel wrist-worn accelerometer data collected in free-living settings for over 24 hours per participant. It includes annotated data from 151 participants, making it significantly larger than existing datasets. We used the first 100 participants as the training set and the remaining 51 as the test set. For each subject, sequences were windowed, and 5% of the data was randomly selected for training and testing. The sample rate was downsampled from 100 Hz to 50 Hz and it includes 10 activity class labels. During the alignment stage, we used a variable window size $w\in[10,500]$, while in the HAR, we fixed $w=500$ with a stride of $250$.
Each dataset includes multiple activity classes, and the proportion of each class in the dataset is shown in Table 10.
Dataset # Classes Classes Proportions (%)
USC-HAD 12
Sleeping, Sitting, Elevator down,
Elevator up, Standing, Jumping,
Walking downstairs, Walking right,
Walking forward, Running forward,
Walking upstairs, Walking left
12.97, 9.06, 6.04, 5.94, 8.6, 3.62,
7.61, 9.81, 13.15, 5.72, 8.22, 9.25
UCI-HAR 6
Standing, Sitting, Laying,
Walking, Walking downstairs,
Walking upstairs
18.69, 17.49, 19.14,
16.68, 13.41, 14.59
PAMAP2 12
Lying, Sitting, Standing,
Ironing, Vacuum cleaning,
Ascending stairs, Descending stairs,
Walking, Nordic walking, Cycling,
Running, Rope jumping
10.25, 9.52, 10.11,
11.82, 9.14, 6.3,
5.67, 12.77, 9.52,
8.42, 3.57, 2.91
MHealth 12
Climbing stairs, Standing still,
Sitting and relaxing, Lying down,
Walking, Waist bends forward,
Frontal elevation of arms,
Knees bending (crouching),
Jogging, Running, Jump front
& back, Cycling
8.91, 8.95, 8.95,
8.95, 8.95, 8.26,
8.7, 8.53, 8.95,
8.95, 2.96, 8.95
CAPTURE-24 10
Sleep, Household-chores, Walking,
Vehicle, Standing, Mixed-activity,
Sitting, Bicycling, Sports,
Manual-work
37.45, 6.5, 6.16,
3.83, 3.25, 3.49,
37.07, 1.03, 0.43, 0.79
Table 10: Dataset classes and Proportions
A.7 Baselines for Task-Aware Tuning Stage